# Text-to-Image Generation Via Energy-Based CLIP

### v1

@@ Abstract
Joint Energy Models (JEMs), while drawing significant research attention, have not been successfully scaled to real-world, high-resolution datasets. We present CLIP-JEM , a novel approach extending JEMs to the multimodal vision-language domain using CLIP, integrating both generative and discriminative objectives. For the generative objective, we introduce an image-text joint-energy function based on Cosine similarity in the CLIP space, training CLIP to assign low energy to real image-caption pairs and high energy otherwise. For the discriminative objective, we employ contrastive adversarial loss, extending the adversarial training objective to the multimodal domain. CLIP-JEM not only generates realistic images from text but also achieves competitive results on the compositionality benchmark, outperforming leading methods with fewer parameters. Additionally, we demonstrate the superior guidance capability of CLIP-JEM by enhancing CLIP-based generative frameworks and converting unconditional diffusion models to text-based ones. Lastly, we show that CLIP-JEM can serve as a more robust evaluation metric for text-to-image generative tasks than CLIP.
@@ 1 Introduction
Energy-based models (EBMs) (LeCun et al. 2006) are a class of models that define a probability distribution over data points using an energy function, where lower energy values correspond to higher probabilities. These models are trained by adjusting the learned function to minimize the energy of observed data points and maximize the energy of synthetic ones, effectively aligning the energy landscape with the true data distribution. Joint Energy Models (JEMs) (Grathwohl et al. 2019) extend EBMs by utilizing a classifier’s logits to also model a joint energy function. JEMs are trained with both discriminative and generative objectives, namely, aiming to classify data points and to model the joint energy function, respectively. However, both EBMs and JEMs face significant scalability challenges. Their training processes can be unstable and computationally intensive, restricting their applicability to smaller datasets and making them unsuitable for real-world, high-resolution image datasets.
@@
Adversarial training (Goodfellow, Shlens, and Szegedy 2015; Madry et al. 2018) is a technique designed to enhance models’ robustness against adversarial examples, which are small, imperceptible perturbations added to inputs to mislead classifiers. By training models to correctly classify these adversarial examples, such training results in perceptually aligned gradients (PAG) (Tsipras et al. 2019). PAG refers to the phenomenon where the model’s input gradients [^1] This gradient is computed as the derivative of the chosen output logit w.r.t. the input image. are semantically meaningful and aligned with human perception, indicating that the features learned by the model are more human-aligned. Recently, the concept of PAG was extended to the multimodal image-text domain with CLIPAG (Ganz and Elad 2024), which applies adversarial training to the visual part of CLIP (Radford et al. 2021). This approach enables text-based generation through pixel-space optimization. However, while CLIPAG can produce good-looking images, the results are often non-realistic and heavily reliant on multiview augmentation. The need for such augmentation suggests that the gradients themselves are not sufficiently informative to generate realistic images from single views. These limitations highlight the need for more advanced techniques to achieve realistic generation using CLIP models.
@@
In this work, we propose CLIP-JEM, a novel approach that extends JEMs to the multimodal vision-language domain using CLIP by combining it with adversarial training. This combination leverages the strengths of both techniques to address their limitations: mitigating the scalability and stability issues of JEMs and enabling high-resolution text-based generation while overcoming the non-realistic outputs typical of CLIPAG. Inspired by unimodal JEMs, we fine-tune CLIP using two objectives: generative and discriminative. For the generative objective, we introduce an image-text energy function based on Cosine similarity in the CLIP space. We train CLIP to assign low-energy values to real image-text pairs and high values to others. More specifically, inspired by EBMs, we utilize the model to draw text-based generated samples and train it to assign these with high energy values. This is done by an iterative pixel-space optimization following the model’s gradients, starting from a random sample. We formulate this as a contrastive loss to align with CLIP. For the discriminative objective, we follow the path set by CLIPAG to define a contrastive adversarial loss. By combining these two objectives, we train the visual encoder of CLIP, resulting in CLIP-JEM, a model with semantically meaningful gradients (fig. 1) capable of generating realistic samples through simple pixel-space optimization (fig. 2).
@@
We establish the effectiveness of CLIP-JEM across three key domains: text-to-image generation, guidance capabilities, and as an evaluation metric. First, CLIP-JEM enables in text-to-image generation through pixel-space optimization, producing realistic images, significantly surpassing CLIPAG by more than $20$ FID points, without applying any augmentation. Despite its relatively small size, this model achieves results competitive with much larger models on the challenging compositionality benchmark, CompBench (Huang et al. 2023). Specifically, it surpasses Stable diffusion v2 (Rombach et al. 2022a) and methods tailored for compositionality (Liu et al. 2023b; Feng et al. 2023). We attribute this to the discriminative nature of the model, enabling it to better align with the provided prompts. Furthermore, CLIP-JEM significantly enhances text-based guiding capabilities. To this end, we illustrate that incorporating CLIP-JEM for guidance converts unconditional diffusion models (Dhariwal and Nichol 2021; Ahn et al. 2024) into text-guided ones with just $25$ diffusion steps. Additionally, replacing CLIP with CLIP-JEM in CLIP-based generative frameworks markedly boosts their performance. Finally, CLIP-JEM proves its utility as an evaluation metric (a.k.a. CLIP-Score) for text-based image editing. It shows robustness to adversarial examples and enhanced sensitivity to image quality compared to the “vanilla” counterpart. This indicates that CLIP-JEM is a more reliable and precise tool for assessing the quality and integrity of generated images. To summarize,
@@ 2.1 Energy-Based Models
EBMs (LeCun et al. 2006) define a probability distribution over data points using an energy function:
@@
where $E_{\theta}(\mathbf{x})$ assigns scalar values to data points, and $Z(\theta)$ is a normalizing constant. Lower energy values indicate higher probabilities. Training such a model with parameters $\theta$ involves minimizing the energy of “positive” data points $\mathbf{x}^{+}\sim p(\mathbf{x})$ and maximizing it for the “negative” samples $\mathbf{x}^{-}\sim p_{\theta}(\mathbf{x})$. The model converges when it cannot distinguish between positive and negative samples. Sampling from $p_{\theta}(\cdot)$ is achieved using Stochastic Gradient Langevin Dynamics (SGLD), starting from a canonical distribution and updating iteratively, using a step-size $\alpha$.
@@ 2.2 Joint-Energy Models
where $f_{\theta}(\mathbf{x})_{y}$ is the logit corresponding with the $y$th class label. Thus, the joint energy function is $E_{\theta}(\mathbf{x},y)=-f_{\theta}(\mathbf{x})_{y}$. Marginalizing over $y$ results in an unconditional distribution $p_{\theta}(\cdot)$. Training JEMs involve with optimizing both a discriminative and a generative objectives. Despite having several merits (e.g., generative capabilities and adversarial robustness), training such models often suffers from instability and even divergence. Despite recent advancements, JEMs perform well for relatively small datasets (mainly SVHN (Netzer et al. 2011), CIFAR (Krizhevsky, Hinton et al. 2009) and CelebA (Liu et al. 2015)) but are not competitive when brought to real-world visual content (Yang, Su, and Ji 2023; Zhu et al. 2021; Yang and Ji 2021). In this work, we aim to extend JEMs into the most challenging setup – text-to-image generation using a CLIP-based model.
@@ 2.3 CLIP for Text-to-Image Generation
CLIP (Radford et al. 2021) is a vision-language model, pretrained to align a massive corpus of image-text pairs. The outstanding performance of CLIP visual and textual encoders has propelled great advancements in various fields. In Large Vision Language Modeling (Li et al. 2022; Zhu et al. 2023; Liu et al. 2023a; Li et al. 2023; Ganz et al. 2023, 2024), CLIP vision encoder serves as the primary visual backbone, leading to unprecedented performance. In text-to-image generation, two main lines of work harness CLIP: (i) Utilizing CLIP image-text alignment to guide the visual results to be aligned with the textual description (Frans, Soros, and Witkowski 2022; Crowson et al. 2022; Patashnik et al. 2021; Gal et al. 2022; Kwon and Ye 2022; Vinker et al. 2022); and (ii) Using CLIP’s text encoder to condition generative models (Kang et al. 2023a; Nichol et al. 2022; Ramesh et al. 2022; Rombach et al. 2022b). Unlike these works, which utilize CLIP along with a generative model, we aim to cast CLIP into an energy-based model, capable of performing text-to-image generation without an additional generative model.
@@ 2.4 Perceptually Aligned Gradients
Adversarially robust models (Carlini and Wagner 2017; Madry et al. 2018) are designed to withstand adversarial attacks (Szegedy et al. 2014; Goodfellow, Shlens, and Szegedy 2015), which are small, imperceptible perturbations aimed at misleading classifiers. It has been observed that such models exhibit a phenomenon known as Perceptually Aligned Gradients (PAG), which is absent in their non-robust counterparts. PAG refers to the model’s input gradients being semantically meaningful and aligned with human perception, indicating that the features learned are more aligned with human vision (Ilyas et al. 2019; Engstrom et al. 2019; Salman et al. 2020). PAG has been harnessed for generative tasks, such as image generation and image-to-image translation (Santurkar et al. 2019), thereby improving state-of-the-art image synthesis results (Ganz and Elad 2022). PAG has also been explored for enhanced robust classification (Blau et al. 2023). Recently, the study of PAG has been extended to the multimodal domain using CLIPAG (Ganz and Elad 2024), which applies adversarial training to the multimodal text-to-image domain using CLIP (Radford et al. 2021). This approach enables text-based generation through pixel-space optimization. However, while CLIPAG can produce good-looking images, the results are often non-realistic and heavily reliant on multiview augmentation. The need for such augmentation suggests that the gradients themselves are not sufficiently informative to generate realistic images from single views. These limitations highlight the need for more advanced techniques for realistic text-to-image generation using CLIP models.
@@ 3 Method
In this work, we aim to extend Joint Energy Models to the challenging text-to-image generation setting using a CLIP model. Our overall framework, illustrated in Figure 3, consists of two main objectives: a contrastive energy loss and an adversarial loss. In section 3.1, we first define the joint image-text energy function, forming a measure of the faithfulness and alignment of the given pair, and elaborate on its training procedure. Next, in section 3.2, we detail the contrastive adversarial loss, extending CLIP’s loss to the adversarial case. Lastly, in section 3.3, we describe the overall training procedure of CLIP-JEM, resulting in a multimodal energy model capable of text-to-image generation.
@@ 3.1 Joint Image-Text Energy Via CLIP
Similar to energy-based model training, adapting CLIP to form an energy measure requires both “positive” and “negative” image-text pairs, denoted as $(\mathbf{I},\mathbf{T})$ and $(\mathbf{\tilde{I}},\mathbf{T})$, respectively. As CLIP is a contrastively-trained model, we utilize such pairs to formulate a contrastive energy objective, illustrated on the upper-left side of Figure 3. This rectangular matrix contains the negative joint-energy values of textual and visual inputs. Specifically, the upper part contains the energy of the “positive” image inputs and the lower one the “negatives”. Given this contrastive energy matrix, we train the model weights of $f_{\theta}^{I}(\cdot)$ using the cross-entropy loss with the main diagonal, marked in blue, as the ground-truth annotations. Namely, the objective is to obtain high values in this main diagonal (low energy, high probability) and low values elsewhere. Focusing on a certain textual input $\mathbf{T}_{k}$, minimizing the proposed contrastive energy loss results in high Cosine similarity values of the “positive” pair $(\mathbf{I}_{k},\mathbf{T}_{k})$ and low ones for the “negative” pair $(\mathbf{\tilde{I}}_{k},\mathbf{T}_{k})$. This formulated loss has two additional advantages: First, minimizing it results in providing low Cosine similarities and hence high energy and low probability for unmatching image-text pairs $(\mathbf{I}_{i},\mathbf{T}_{j})$ for $i\neq j$, enhancing the model’s discriminative ability. Second, it enables the utilization of different realizations of negative samples per each text prompt, which we empirically find to stabilize the training.
@@
The remaining question is how to obtain the “negative” samples. Similar to the methodology in the JEM line of work, we craft such samples in an iterative process based on the model’s gradients, starting from a simple canonical distribution (e.g., a Unifrom distribution). Current JEM works utilize SGLD (Welling and Teh 2011), which requires hundreds of iterations to obtain good samples. To mitigate the computation overhead involved in generating such samples, such works utilize a replay buffer. This data structure stores thousands of generated “negative” samples and updates them throughout the training process, enabling a relatively small number of SGLD steps. However, in the image-text context, employing such a solution is not feasible, as, unlike multiclass datasets such as CIFAR-10, which have a fixed amount of classes, the text-to-image setting is open-vocabulary in nature with practically infinite number of possible captions. Thus, in our setting, maintaining a replay buffer and updating its samples is not a practical solution, as the probability of having the exact two captions in a dataset is very low, eliminating the usefulness of such a structure. To enable fast “negative” sampling, we propose to avoid employing SGLD and utilize a momentum-based optimization with an adaptive learning rate instead. In practice, we use AdamW optimizer (Loshchilov and Hutter 2017) to update the image in the iterative process, which enables drawing samples within a relatively small number of steps. Overall, we first draw the initial sample from a canonical uniform distribution, $\tilde{I}^{t=0}$. Next, we compute the Cosine similarity between the visual and textual encodings and calculate its input-gradients. Lastly, we update the visual inputs accordingly and repeat the process for $T_{\text{JEM}}$ iterations. Notably, we calculate the gradients on a slightly noisy version of the visual input, introducing randomness to the sampling mechanism. In the supplementary materials, we discuss the connection of our sampling procedure to SGLD.

### v2

@@ Abstract
Joint Energy Models (JEMs), while drawing significant research attention, have not been successfully scaled to real-world, high-resolution datasets. We present CLIP-JEM , a novel approach extending JEMs to the multimodal vision-language domain using CLIP, integrating both generative and discriminative objectives. For the generative one, we introduce an image-text joint-energy function based on Cosine similarity in the CLIP space, training CLIP to assign low energy to real image-caption pairs and high energy otherwise. For the discriminative one, we employ contrastive adversarial loss, extending the adversarial training objective to the multimodal domain. CLIP-JEM not only generates realistic images from text but also achieves competitive results on the compositionality benchmark, outperforming leading methods with fewer parameters. Additionally, we demonstrate the superior guidance capability of CLIP-JEM by enhancing CLIP-based generative frameworks and converting unconditional diffusion models to text-based ones. Lastly, we show that our model can serve as a more robust evaluation metric for text-to-image generative tasks than CLIP.
@@ 1 Introduction
Energy-based models (EBMs) (LeCun et al., 2006) are a class of models that define a probability distribution over data points using an energy function, where lower energy values correspond to higher probabilities. These models are trained by adjusting the learned function to minimize the energy of observed data points and maximize the energy of synthetic ones, effectively aligning the energy landscape with the true data distribution. Joint Energy Models (JEMs) (Grathwohl et al., 2019) extend EBMs by utilizing a classifier’s logits to also model a joint energy function. JEMs are trained with both discriminative and generative objectives, namely, aiming to classify data points and to model the joint energy function, respectively. However, both EBMs and JEMs face significant scalability challenges. Their training processes can be unstable and computationally intensive, restricting their applicability to smaller datasets and making them unsuitable for real-world, high-resolution image datasets.
@@
Adversarial training (Goodfellow et al., 2015; Madry et al., 2018) is a technique designed to enhance models’ robustness against adversarial examples, which are small, imperceptible perturbations added to inputs to mislead classifiers. By training models to correctly classify these adversarial examples, such training results in perceptually aligned gradients (PAG) (Tsipras et al., 2019). PAG refers to the phenomenon where the model’s input gradients [^1] This gradient is computed as the derivative of the chosen output logit w.r.t. the input image. are semantically meaningful and aligned with human perception, indicating that the features learned by the model are more human-aligned. Recently, the concept of PAG was extended to the multimodal image-text domain with CLIPAG (Ganz & Elad, 2024), which applies adversarial training to the visual part of CLIP (Radford et al., 2021). This approach enables text-based generation through pixel-space optimization. However, while CLIPAG can produce good-looking images, the results are often non-realistic and heavily reliant on multiview augmentation. The need for such augmentation suggests that the gradients themselves are not sufficiently informative to generate realistic images from single views. These limitations highlight the need for more advanced techniques to achieve realistic generation using CLIP models.
@@
In this work, we propose CLIP-JEM, a novel approach that extends JEMs to the multimodal vision-language domain using CLIP by combining it with adversarial training. This combination leverages the strengths of both techniques to address their limitations: mitigating the scalability and stability issues of JEMs and enabling high-resolution text-based generation while overcoming the non-realistic outputs typical of CLIPAG. Inspired by unimodal JEMs, we fine-tune CLIP using two objectives: generative and discriminative. For the generative objective, we introduce an image-text energy function based on Cosine similarity in the CLIP space. We train CLIP to assign low-energy values to real image-text pairs and high values to others. More specifically, inspired by EBMs, we utilize the model to draw text-based generated samples and train it to assign these with high energy values. This is done by an iterative pixel-space optimization following the model’s gradients, starting from a random sample. We formulate this as a contrastive loss to align with CLIP. For the discriminative objective, we follow the path set by CLIPAG to define a contrastive adversarial loss. By combining these two objectives, we train the visual encoder of CLIP, resulting in CLIP-JEM, a model with semantically meaningful gradients (fig.˜1) capable of generating realistic samples through simple pixel-space optimization (fig.˜2).
@@
We establish the effectiveness of CLIP-JEM across three key domains: text-to-image generation, guidance capabilities, and as an evaluation metric. First, CLIP-JEM enables in text-to-image generation through pixel-space optimization, producing realistic images, significantly surpassing CLIPAG by more than 202020 FID points, without applying any augmentation. Despite its relatively small size, this model achieves results competitive with much larger models on the challenging compositionality benchmark, CompBench (Huang et al., 2023). Specifically, it surpasses Stable diffusion v2 (Rombach et al., 2022a) and methods tailored for compositionality (Liu et al., 2023b; Feng et al., 2023). We attribute this to the discriminative nature of the model, enabling it to better align with the provided prompts. Furthermore, CLIP-JEM significantly enhances text-based guiding capabilities. To this end, we illustrate that incorporating CLIP-JEM for guidance converts unconditional diffusion models (Dhariwal & Nichol, 2021; Ahn et al., 2024) into text-guided ones with just 252525 diffusion steps. Additionally, replacing CLIP with CLIP-JEM in CLIP-based generative frameworks markedly boosts their performance. Finally, CLIP-JEM proves its utility as an evaluation metric (a.k.a. CLIP-Score) for text-based image editing. It shows robustness to adversarial examples and enhanced sensitivity to image quality compared to the “vanilla” counterpart. This indicates that CLIP-JEM is a more reliable and precise tool for assessing the quality and integrity of generated images. To summarize,
@@ 2.1 Energy-Based Models
EBMs (LeCun et al., 2006) define a probability distribution over data points using an energy function:
@@
where $E_{\theta}(\mathbf{x})$ assigns scalar values to data points, and $Z(\theta)$ is a normalizing constant. Lower energy values indicate higher probabilities. Training a model with parameters $\theta$ involves minimizing the energy assigned to “positive” data samples $\mathbf{x}^{+}\sim p(\mathbf{x})$ and maximizing the energy of “negative” samples $\mathbf{x}^{-}\sim p_{\theta}(\mathbf{x})$. In this framework, the model’s ability to “distinguish” between samples is operationalized by assigning lower energy (i.e., higher likelihood) to samples drawn from the data distribution, and higher energy (i.e., lower likelihood) to those drawn from the model distribution. Convergence is achieved when the model can no longer reliably separate positive and negative samples based on their energy values. To sample from $p_{\theta}(\cdot)$, we employ Stochastic Gradient Langevin Dynamics (SGLD), which begins from a predefined initial distribution and iteratively updates the samples using a step size $\alpha$.
@@ 2.2 Joint-Energy Models
where $f_{\theta}(\mathbf{x})_{y}$ is the logit corresponding with the $y$th class label. Thus, the joint energy function is $E_{\theta}(\mathbf{x},y)=-f_{\theta}(\mathbf{x})_{y}$. Marginalizing over $y$ results in an unconditional distribution $p_{\theta}(\cdot)$. Training JEMs involve with optimizing both a discriminative and a generative objectives. Despite having several merits (e.g., generative capabilities and adversarial robustness), training such models often suffers from instability and even divergence. Despite recent advancements, JEMs perform well for relatively small datasets (mainly SVHN (Netzer et al., 2011), CIFAR (Krizhevsky et al., 2009) and CelebA (Liu et al., 2015)) but are not competitive when brought to real-world visual content (Yang et al., 2023; Zhu et al., 2021; Yang & Ji, 2021). In this work, we aim to extend JEMs into the most challenging setup – text-to-image generation using a CLIP-based model.
@@ 2.3 CLIP for Text-to-Image Generation
CLIP (Radford et al., 2021) is a vision-language model, pretrained to align a massive corpus of image-text pairs. The outstanding performance of CLIP visual and textual encoders has propelled great advancements in various fields. In Large Vision Language Modeling (Li et al., 2022; Zhu et al., 2023; Liu et al., 2023a; Li et al., 2023; Ganz et al., 2023; 2024), CLIP vision encoder serves as the primary visual backbone, leading to unprecedented performance. In text-to-image generation, two main lines of work harness CLIP: (i) Utilizing CLIP image-text alignment to guide the visual results to be aligned with the textual description (Frans et al., 2022; Crowson et al., 2022; Patashnik et al., 2021; Gal et al., 2022; Kwon & Ye, 2022; Vinker et al., 2022); and (ii) Using CLIP’s text encoder to condition generative models (Kang et al., 2023a; Nichol et al., 2022; Ramesh et al., 2022; Rombach et al., 2022b). Unlike these works, which utilize CLIP along with a generative model, we aim to cast CLIP into an energy-based model, capable of performing text-to-image generation without an additional generative model.
@@ 2.4 Perceptually Aligned Gradients
Adversarially robust models (Carlini & Wagner, 2017; Madry et al., 2018) are designed to withstand adversarial attacks (Szegedy et al., 2014; Goodfellow et al., 2015), which are small, imperceptible perturbations aimed at misleading classifiers. It has been observed that such models exhibit a phenomenon known as Perceptually Aligned Gradients (PAG), which is absent in their non-robust counterparts. PAG refers to the model’s input gradients being semantically meaningful and aligned with human perception, indicating that the features learned are more aligned with human vision (Ilyas et al., 2019; Engstrom et al., 2019; Salman et al., 2020). PAG has been harnessed for generative tasks, such as image generation and image-to-image translation (Santurkar et al., 2019), thereby improving state-of-the-art image synthesis results (Ganz & Elad, 2022). PAG has also been explored for enhanced robust classification (Blau et al., 2023). Recently, the study of PAG has been extended to the multimodal domain using CLIPAG (Ganz & Elad, 2024), which applies adversarial training to the multimodal text-to-image domain using CLIP (Radford et al., 2021). This approach enables text-based generation through pixel-space optimization. However, while CLIPAG can produce good-looking images, the results are often non-realistic and heavily reliant on multiview augmentation. The need for such augmentation suggests that the gradients themselves are not sufficiently informative to generate realistic images from single views. These limitations highlight the need for more advanced techniques for realistic text-to-image generation using CLIP models.
@@ 3 Method
In this work, we aim to extend Joint Energy Models to the challenging text-to-image generation setting using a CLIP model. Our overall framework, illustrated in Figure˜3, consists of two main objectives: a contrastive energy loss and an adversarial loss. In section˜3.1, we first define the joint image-text energy function, forming a measure of the faithfulness and alignment of the given pair, and elaborate on its training procedure. Next, in section˜3.2, we detail the contrastive adversarial loss, extending CLIP’s loss to the adversarial case. Lastly, in section˜3.3, we describe the overall training procedure of CLIP-JEM, resulting in a multimodal energy model capable of text-to-image generation.
@@ 3.1 Joint Image-Text Energy Via CLIP
Similar to energy-based model training, adapting CLIP to form an energy measure requires both “positive” and “negative” image-text pairs, denoted as $(\mathbf{I},\mathbf{T})$ and $(\mathbf{\tilde{I}},\mathbf{T})$, respectively. As CLIP is a contrastively-trained model, we utilize such pairs to formulate a contrastive energy objective, illustrated on the upper-left side of Figure˜3. This rectangular matrix contains the negative joint-energy values of textual and visual inputs. Specifically, the upper part contains the energy of the “positive” image inputs and the lower one the “negatives”. Given this contrastive energy matrix, we train the model weights of $f_{\theta}^{I}(\cdot)$ using the cross-entropy loss with the main diagonal, marked in blue, as the ground-truth annotations. Namely, the objective is to obtain high values in this main diagonal (low energy, high probability) and low values elsewhere. Focusing on a certain textual input $\mathbf{T}_{k}$, minimizing the proposed contrastive energy loss results in high Cosine similarity values of the “positive” pair $(\mathbf{I}_{k},\mathbf{T}_{k})$ and low ones for the “negative” pair $(\mathbf{\tilde{I}}_{k},\mathbf{T}_{k})$. This formulated loss has two additional advantages: First, minimizing it results in providing low Cosine similarities and hence high energy and low probability for unmatching image-text pairs $(\mathbf{I}_{i},\mathbf{T}_{j})$ for $i\neq j$, enhancing the model’s discriminative ability. Second, it enables the utilization of different realizations of negative samples per each text prompt, which we empirically find to stabilize the training.
@@
The remaining question is how to obtain the “negative” samples. Similar to the methodology in the JEM line of work, we craft such samples in an iterative process based on the model’s gradients, starting from a simple canonical distribution (e.g., a Unifrom distribution). Current JEM works utilize SGLD (Welling & Teh, 2011), which requires hundreds of iterations to obtain good samples. To mitigate the computation overhead involved in generating such samples, such works utilize a replay buffer. This data structure stores thousands of generated “negative” samples and updates them throughout the training process, enabling a relatively small number of SGLD steps. However, in the image-text context, employing such a solution is not feasible, as, unlike multiclass datasets such as CIFAR-10, which have a fixed amount of classes, the text-to-image setting is open-vocabulary in nature with practically infinite number of possible captions. Thus, in our setting, maintaining a replay buffer and updating its samples is not a practical solution, as the probability of having the exact two captions in a dataset is very low, eliminating the usefulness of such a structure. To enable fast “negative” sampling, we propose to avoid employing SGLD and utilize a momentum-based optimization with an adaptive learning rate instead. In practice, we use AdamW optimizer (Loshchilov & Hutter, 2017) to update the image in the iterative process, which enables drawing samples within a relatively small number of steps. Overall, we first draw the initial sample from a canonical uniform distribution, $\tilde{I}^{t=0}$. Next, we compute the Cosine similarity between the visual and textual encodings and calculate its input-gradients. Lastly, we update the visual inputs accordingly and repeat the process for $T_{\text{JEM}}$ iterations. Notably, we calculate the gradients on a slightly noisy version of the visual input, introducing randomness to the sampling mechanism. In the supplementary materials, we discuss the connection of our sampling procedure to SGLD.